Circular No. 8591

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005dp IN NGC 5630 S. Nakano, Sumoto, Japan, reports the discovery by Koichi Itagaki

(Teppo-cho, Yamagata, Japan) of an apparent supernova (red mag 16.0) on unfiltered CCD images taken around Aug. 27.48 and 29.42 UT with a 0.60-m f/5.7 reflector. The new object is located at α = 14^{h}27^{m}36^{s}.62, δ = +41^{o}15′15′′.0 (equinox 2000.0), which is 1′′ west and 14′′ south of the center of NGC 5630. Nothing was visible at this location on images taken on 2002 May 2 and 2005 Aug. 5 (limiting mag 19.0).

VARIABLE STAR NEAR M27 W. Renz, Karlsruhe, Germany, reports the CCD discovery of a variable

star situated in the outer portion of M27 by J. Hanisch (Gescher, Germany) on unfiltered CCD images taken on ∼ Aug. 18.0 UT (at mag ∼ 15.8; position α = 19^{h}59^{m}51^{s}.29, δ = +22^{o}42′32′′.3, equinox 2000.0) and independently by H.-G. Lindberg (Skultuna, Sweden) on an image taken on Aug. 18.9. Renz also forwards position end figures 51^{s}.283, 32′′.13 obtained by D. Boyd (Oxfordshire, U.K., 0.35-m reflector) from CCD frames taken on Aug. 21.92. CCD magnitudes for the variable star (reported in part by E. Waagen, AAVSO): Aug. 16.937, 15.6 (N. Quinn, West Sussex, U.K.; prediscovery; unfiltered); 21.916, V = 16.17 (Boyd); 21.917, I = 16.00 (Boyd); 21.935, V = 16.69 (D. Rodriguez, Madrid, Spain); 22.998, V = 16.16 (Boyd); 22.997, I_{c} = 15.82 (Boyd); 26.194–26.202, R = 15.6 (J. E. McGaha, Tucson, AZ); 27.241, V = 15.6 (McGaha); 27.245, R = 15.8 (McGaha); 27.249, B = 14.9 (McGaha). A. Henden, AAVSO, writes that a likely progenitor with R_{c} ∼ 22.5 has been identified on co-added U.S. Naval Observatory images of M27 taken in the early 1990s (limiting mag R_{c} ∼ 26, but a difficult measurement with the variable background and nearby brighter stars) — giving an amplitude of ∼ 7 magnitudes, typical for a WZ-Sge-type cataclysmic variable. Renz and Henden add that ‘superhump’ variations in brightness of ∼ 0.3 mag with a period of ∼ 82 min have been detected in photometric observations by T. Vanmunster and B. Martin.

COMETS 169P/2002 EX_{12} (NEAT), 170P/2005 M1 (CHRISTENSEN)

Comet P/2002 EX_{12} (cf. IAUC 8578; found in images dating back to 1988) has been given the permanent numbering 169P, while comet P/2005 M1 (cf. IAUC 8547; found in images dating back to 1997) has been assigned the number 170P (MPC 54661).

2005 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green